OpenStudy (anonymous):

In a sequence of four positive numbers, the first three are in geometric progression and the last three are in arithmetic progression. The first number is 12 and the last number is 452. The sum of the two middle numbers can be written as ab where a and b are coprime positive integers. Find a+b.

OpenStudy (cwrw238):

the first 3 numbers are 12 12r and 12r^2 where r = common ratio of the G P last 3 are 12r , 12r + d and 12r + 2d where d = common difference of the A P 12r + 12d = 45

OpenStudy (cwrw238):

* 12r + 2d = 452

OpenStudy (amistre64):

my idea: 12, 12+k, 12+2k (12+k)r = 12+2k 12, 12+k, 12+2k, (12+k)r^2 = 452 (12+k)r^2 = 452

OpenStudy (amistre64):

... lol, i got that backwards dont i

OpenStudy (cwrw238):

yea lol

OpenStudy (cwrw238):

it comes down to solving system of 2 equations, i think 12r + 2d = 452 and 12r + d = 12r^2 ( the third term)

OpenStudy (anonymous):

12, 12r, 12r^2, 452. since the last three terms 12r, 12r^2 and 452 are in arithmetic progression, 12r^2=(12r+452)/2

OpenStudy (cwrw238):

from first equation 2d = 452 - 12r d = 226 - r plug this into second equation to get 12r + 226 - r = 12r^2 12r^2 - 11r - 226 = 0

OpenStudy (cwrw238):

ooohh - i see i have have made a mistake d = 226 - 6r

OpenStudy (anonymous):

@cwrw238 yea..right. even i'm getting the same thing! :)

OpenStudy (cwrw238):

lol

OpenStudy (cwrw238):

common ratio r satisfies the equation 12r^2 - 6r - 226 = 0

OpenStudy (cwrw238):

solving this gives r = 4.569 or -4.0969

OpenStudy (cwrw238):

we can ignore the negative root because the 4 numbers are all postive d = (452 - 12(4.5969) ) / 2 = 198.4186 this makes the sequence 12, 55.1628, 253.5814, 452